ROTATIONAL SPECTRA OF THE H_{2}-HCN COMPLEX OBSERVED BY FOURIER-TRANSFORM MI-CROWAVE SPECTROSCOPY. THE INTERNAL ROTATION OF H_{2} IN THE H_{2}-HCN COMPLEX.

M. ISHIGURO, K. HARADA, \underline{KEIICHI TANAKA} AND T. TANAKA, Department of Chemistry, Faculty of Science, Kyushu University, Fukuoka, 812-8581 Japan; Y. SUMIYOSHI AND Y. ENDO, Department of Pure and Applied Sciences, Graduate School of Arts and Sciences, The University of Tokyo, Komaba, Meguroku, Tokyo, 153-8902, Japan.

Fourier-Transform Microwave (FTMW) spectroscopy has been applied to the observation of the J = 1 − 0 rotational lines of the H_{2}-HCN and H_{2}-DCN complexes for the ortho- and para-H_{2} species. The rotational constants derived for the HCN/DCN complexes support definitely that H_{2} is attached to the N end of HCN for the ortho-H_{2} species, while H_{2} is attached to the H end of HCN for the para-H_{2} species, consistent with the recent vibrational spectra in the He-droplet ^{a}. The nuclear quadrupole interaction constants of the nitrogen nucleus (I_{N} = 1) show that the HCN/DCN part executes a floppy motion with a large mean amplitude. The splitting due to the magnetic hyperfine interaction of the H_{2} part was observed for ortho-H_{2}-HCN (j = 1 for H_{2} internal rotation) thanks to the high resolution of FTMW spectroscopy, but not for para-H_{2}-HCN (j = 0 for H_{2} internal rotation). Because of Σ symmetry of the ortho-H_{2}-HCN complex in the ground state, nuclear spin-nuclear spin interaction contributes to the hydrogen hyperfine interaction, but not the nuclear spin-rotation intraction. The magnetic interaction constant d_{H} for ortho-H_{2}-HCN was determined to be 54.6(38) kHz. The value corresponds to 94.7% of that of free hydrogen molecule 57.671(24) kHz ^{b}, indicating almost free rotation of the H_{2} part in the complex. The spectrum of ortho-H_{2}-DCN shows also the splitting due to the magnatic hyperfine interaction of H_{2}, but the hyperfine partern is more complicated because of the nuclear quadrupole interaction of the D atom (I_{D} = 1). The magnetic interaction constant for ortho-H_{2}-DCN indicates that H_{2} is rotating almost freely in the complex as well.

aD. T .Moore, M. Ishiguro, and R. E. Miller, J. Chem. P hys. \underline{115}, 5144 (2001)

bH. G. Kolsky and N. F. Ramsey, P hys. Rev. \underline{87}, 395 (1952)